OpenStudy (credmond):

Find the derivative of the function defined as follows: y= -8e^tanx

hartnn (hartnn):

do you know chain rule ??

OpenStudy (credmond):

mhm

hartnn (hartnn):

so what answer did you get for that ?

hartnn (hartnn):

\(\Large y'= (-8e^{\tan x})' = -8 (e^{\tan x})(\tan x)' = ...\)

OpenStudy (credmond):

sooo... e^tanx = sin(x)/cos(x)^2 right?

OpenStudy (isaiah.feynman):

Nope.

hartnn (hartnn):

there is no need to convert tan to sin/cos here do you know the derivative of tan x ??

OpenStudy (credmond):

ohhh. okay

OpenStudy (credmond):

it's (sec x) ^2 right?

OpenStudy (isaiah.feynman):

Yes.

hartnn (hartnn):

so, whats your answer ?

OpenStudy (credmond):

so is it -16(sec x)(e^tanx) ?

hartnn (hartnn):

\(\Large y'= (-8e^{\tan x})' = -8 (e^{\tan x})(\tan x)' = -8 (e^{\tan x}) \sec^2 x \)

OpenStudy (isaiah.feynman):